Hey, could you please help me with this word problem. I can't figure it out. A car radiator is filled with 5L of a 25% antifreeze solution. How many liters must be drawn off and replaced by a 75% antifreeze solution to leave the radiator filled with a 55% antifreeze solution?
Thank you.

Answer by Paul(988)

5(25)+75(5-x)=55(5)
125+375-75x=275
-75x=-225
x=3

Hence, about 3L of solution must be replaced.
Paul.